# Non-commutative Iwasawa theory of abelian varieties over global function fields

### v1

@@ Abstract.
Let $A$ be an abelian variety defined over a global function field $F$. We investigate the structure of the $p$-primary Selmer group $\mathrm{Sel}(A/F_{\infty})$ for any prime number $p$ distinct from the characteristic of $F$, over $p$-adic Lie extensions $F_{\infty}$ of $F$ which contain the cyclotomic $\mathbb{Z}_{p}$-extension $F^{\mathrm{cyc}}$. In particular, we prove that the Pontryagin dual of the Selmer group $\mathrm{Sel}(A/F^{\mathrm{cyc}})$ is a torsion $\mathbb{Z}_{p}[[\mathrm{Gal}(F^{\mathrm{cyc}}/F)]]$-module with trivial $\mu$-invariant, and we establish the $\mathfrak{M}_{H}(G)$-conjecture of Coates–Fukaya–Kato–Sujatha–Venjakob for $A/F_{\infty}$.
In view of the validity of the $\mathfrak{M}_{H}(G)$-conjecture, it therefore makes sense to speak of the characteristic element (in the sense of Coates et al.) attached to the Pontryagin dual of $\mathrm{Sel}(A/F_{\infty})$. We relate the order of vanishing of the characteristic elements, evaluated at Artin representations, to the corank of the Selmer group of the corresponding twist of $A$ over the base field $F$. Combining this with the deep results of Tate, Milne and Kato–Trihan, we show that the order of vanishing of the characteristic elements is equal to the order of vanishing of the $L$-function of $A/F$ at $s=1$ under appropriate assumptions.
Finally, we relate the generalised Euler characteristic of $\mathrm{Sel}(A/F_{\infty})$ to the Euler characteristic of $\mathrm{Sel}(A/F^{\mathrm{cyc}})$. This is a natural analogue of Zerbes’ result in the number field context and generalises previous results of Sechi and Valentino in the function field context.
@@
Elliptic curves, Iwasawa theory, Euler characteristics
@@
The second author is supported by the ANR project Coloss and the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No. 101026826.
@@ 1. Introduction
Non-commutative Iwasawa theory has emerged as a powerful framework for understanding deep arithmetic properties over number fields contained in a $p$-adic Lie extension and their precise relationship to special values of complex $L$-functions. The aim of the present paper is to study non-commutative Iwasawa theory over global function fields. To motivate our discussion, we begin by reviewing the number field context as developed by Coates, Fukaya, Kato, Sujatha, and Venjakob in [8] (see also [3, 4, 38]). In the number field case, $p$ will denote an odd prime. Given a compact $p$-adic Lie group $\mathcal{G}$, the Iwasawa algebra $\Lambda(\mathcal{G})$ of $\mathcal{G}$ is defined to be the completed group ring
@@
where $\mathcal{U}$ runs over all open normal subgroups of $\mathcal{G}$. Let $E$ be an elliptic curve defined over a number field $F$ with good ordinary reduction at all primes of $F$ above $p$. Denote by $F_{\infty}$ an admissible $p$-adic Lie extension of $F$, that is, a Galois extension of $F$ such that $F^{\mathrm{cyc}}\subset F_{\infty}$ and whose Galois group $G={\mathrm{Gal}}(F_{\infty}/F)$ is a $p$-adic Lie group of positive dimension. Here $F^{\mathrm{cyc}}$ denotes the cyclotomic $\mathbb{Z}_{p}$-extension of $F$. Write $\Gamma={\mathrm{Gal}}(F^{\mathrm{cyc}}/F)$, $H={\mathrm{Gal}}(F_{\infty}/F^{\mathrm{cyc}})$, and set
$X(E/F_{\infty})={\mathrm{Hom}}_{\rm cts}({\mathrm{Sel}}(E/F_{\infty}),{\mathbb{Q}}_{p}/{\mathbb{Z}}_{p})$
for the Pontryagin dual of the $p$-primary Selmer group ${\mathrm{Sel}}(E/F_{\infty})$, endowed with its natural left $\Lambda(G)$-module structure. In the case where $F_{\infty}=F^{\mathrm{cyc}}$, it is well-known that $X(E/F^{\mathrm{cyc}})$ is a finitely generated $\Lambda(\Gamma)$-module. It is generally expected that $X(E/F^{\mathrm{cyc}})$ enjoys nice module-theoretical properties, as illustrated by the following conjecture of Mazur [24]:
@@
This has been verified by Mazur himself in the case where ${\mathrm{Sel}}(E/F)$ is finite. The best evidence in support of this conjecture is a deep result of Kato [17], where the torsionness assertion is established when $E$ is defined over ${\mathbb{Q}}$ and $F$ is an abelian extension of ${\mathbb{Q}}$. Taking the structure theory of $\Lambda(\Gamma)$-module into account, Mazur’s conjecture is equivalent to saying that $X(E/F^{\mathrm{cyc}})/X(E/F^{\mathrm{cyc}})(p)$ is finitely generated over ${\mathbb{Z}}_{p}$, where $X(E/F^{\mathrm{cyc}})(p)$ denotes the $p$-primary submodule of $X(E/F^{\mathrm{cyc}})$. Motivated by this observation, Coates et al. introduce the category ${\mathfrak{M}}_{H}(G)$ which consists of all finitely generated $\Lambda(G)$-modules $M$ whose quotient $M/M(p)$ by the $p$-primary submodule $M(p)$ is a finitely generated $\Lambda(H)$-module. They then conjectured that ${\mathfrak{M}}_{H}(G)$ contains all the torsion $\Lambda(G)$-modules which are of interest in arithmetic applications. More precisely, they formulated (see Section 5 and Conjecture 5.1 of [8]):
@@
This conjecture is crucial as it allows one to define a characteristic element for $X(E/F_{\infty})$. This is essential for formulating the main conjectures of non-commutative Iwasawa theory (see [8, 4]). Coates et al. further demonstrated in their paper how these characteristic elements relate to the Euler characteristics of $X(E/F_{\infty})$ via the theory of Akashi series introduced in [9]. This latter aspect has been further explored by Zerbes in a series of papers [39, 40, 41], where she related the characteristic elements to the generalised Euler characteristics of $X(E/F_{\infty})$. Here, the notion of the generalised Euler characteristics is an extension of the usual Euler characteristics, whose idea emerged in [9], and was later developed and systematically studied by Zerbes in [40].
@@
We now shift our focus to the function field setting, the central theme of our paper. Henceforth, $p$ denotes a fixed prime, where we allow $p=2$. Throughout this paper, $F$ will always denote a global function field. In other words, $F$ is a finite extension of the field $k(T)$ of rational functions in one variable over a finite constant field $k$. We always assume that the characteristic $\mathrm{char}(F)$ of $F$ is distinct from our fixed prime $p$. We fix a separable closure $\bar{F}$ of $F$. The cyclotomic ${\mathbb{Z}}_{p}$-extension $F^{\rm cyc}$ of $F$ is obtained by adjoining to $F$ the unique ${\mathbb{Z}}_{p}$-constant field extension of $k$. Thus, writing $\mu_{p^{\infty}}$ for the group of $p$-power roots of unity in an algebraic closure $\bar{k}$ of $k$, $F^{\rm cyc}$ is the unique ${\mathbb{Z}}_{p}$-extension of $F$ contained in $F(\mu_{p^{\infty}})$. Following the analogy with the number field setting, we define an admissible $p$-adic Lie extension of $F$ as an extension $F_{\infty}$ satisfying the following conditions:

### v2

@@ Abstract.
Let $A$ be an abelian variety defined over a global function field $F$, and let $p$ be a prime distinct from the characteristic of $F$. Let $F_{\infty}$ be a $p$-adic Lie extension of $F$ that contains the cyclotomic $\mathbb{Z}_{p}$-extension $F^{\mathrm{cyc}}$ of $F$. In this paper, we investigate the structure of the $p$-primary Selmer group $\mathrm{Sel}(A/F_{\infty})$ of $A$ over $F_{\infty}$. We prove the $\mathfrak{M}_{H}(G)$-conjecture for $A/F_{\infty}$. Furthermore, we show that both the $\mu$-invariant of the Pontryagin dual of the Selmer group $\mathrm{Sel}(A/F^{\mathrm{cyc}})$ and the generalised $\mu$-invariant of the Pontryagin dual of the Selmer group $\mathrm{Sel}(A/F_{\infty})$ are zero, therby proving Mazur’s conjecture for $A/F$. We then relate the order of vanishing of the characteristic elements, evaluated at Artin representations, to the corank of the Selmer group of the corresponding twist of $A$ over the base field $F$. Assuming the finiteness of the Tate–Shafarevich group, we establish that this corank equals the order of vanishing of the $L$-function of $A/F$ at $s=1$. Finally, we extend a theorem of Sechi—originally proved for elliptic curves without complex multiplication—to abelian varieties over global function fields. This is achieved by adapting the notion of generalised Euler characteristic, introduced by Zerbes for elliptic curves over number fields. This new invariant allows us, via Akashi series, to relate the generalised Euler characteristic of $\mathrm{Sel}(A/F_{\infty})$ to the Euler characteristic of $\mathrm{Sel}(A/F^{\mathrm{cyc}})$.
@@
Abelian varieties, Iwasawa theory, global function fields, Generalised Euler characteristics
@@ 1. Introduction
Non-commutative Iwasawa theory has emerged as a powerful framework for understanding deep arithmetic properties over number fields contained in a $p$-adic Lie extension and their precise relationship with special values of complex $L$-functions. The aim of the present paper is to extend the study of non-commutative Iwasawa theory to the setting of global function fields. To motivate our discussion, we begin by reviewing the number field context as developed by Coates, Fukaya, Kato, Sujatha, and Venjakob in [9] (see also [3, 4, 47]). In the number field setting, let $p$ denote an odd prime. Given a compact $p$-adic Lie group $\mathcal{G}$, the Iwasawa algebra $\Lambda(\mathcal{G})$ of $\mathcal{G}$ is defined as the completed group ring
@@
where the inverse limit is taken over all open normal subgroups $\mathcal{U}$ of $\mathcal{G}$. Let $E$ be an elliptic curve defined over a number field $F$, with good ordinary reduction at all primes of $F$ lying above $p$. Let $F_{\infty}$ be a $p$-adic Lie extension of $F$ such that the cyclotomic ${\mathbb{Z}}_{p}$-extension $F^{\mathrm{cyc}}$ of $F$ is contained in $F_{\infty}$, and the Galois group $G={\mathrm{Gal}}(F_{\infty}/F)$ is a $p$-adic Lie group of positive dimension. Write $\Gamma={\mathrm{Gal}}(F^{\mathrm{cyc}}/F)$, $H={\mathrm{Gal}}(F_{\infty}/F^{\mathrm{cyc}})$, and define
$X(E/F_{\infty}):={\mathrm{Hom}}_{\rm cts}({\mathrm{Sel}}(E/F_{\infty}),{\mathbb{Q}}_{p}/{\mathbb{Z}}_{p}),$
the Pontryagin dual of the $p$-primary Selmer group ${\mathrm{Sel}}(E/F_{\infty})$, endowed with its natural left $\Lambda(G)$-module structure. In the case where $F_{\infty}=F^{\mathrm{cyc}}$, it is well known that $X(E/F^{\mathrm{cyc}})$ is a finitely generated $\Lambda(\Gamma)$-module. It is generally expected that $X(E/F^{\mathrm{cyc}})$ enjoys favorable module-theoretic properties, as reflected in the following conjecture of Mazur [31]:
@@
Mazur verified this conjecture in the case where ${\mathrm{Sel}}(E/F)$ is finite. The strongest evidence supporting this conjecture is a deep result of Kato [21], who established the torsion property when $E$ is defined over $\mathbb{Q}$} and $F$ is an abelian extension of $\mathbb{Q}$}. Taking into account the structure theory of $\Lambda(\Gamma)$-modules, Mazur’s conjecture is equivalent to asserting that the quotient $X(E/F^{\mathrm{cyc}})/X(E/F^{\mathrm{cyc}})(p)$ is finitely generated over ${\mathbb{Z}}_{p}$, where $X(E/F^{\mathrm{cyc}})(p)$ denotes the ${\mathbb{Z}}_{p}$-torsion submodule of $X(E/F^{\mathrm{cyc}})$. Motivated by this observation, Coates and collaborators introduced the category ${\mathfrak{M}}_{H}(G)$, consisting of all finitely generated $\Lambda(G)$-modules $M$ such that the quotient $M/M(p)$ by its ${\mathbb{Z}}_{p}$-torsion submodule $M(p)$ is a finitely generated $\Lambda(H)$-module. They then conjectured that this category contains all the torsion $\Lambda(G)$-modules of arithmetic significance. More precisely, they formulated (see Section 5 and Conjecture 5.1 of [9]):
@@
This conjecture is crucial, as it enables the definition of a characteristic element for $X(E/F_{\infty})$, which is essential for formulating the main conjectures in non-commutative Iwasawa theory (see [9, 4]). Coates et al. further demonstrated how these characteristic elements relate to the Euler characteristics of $X(E/F_{\infty})$ via the theory of Akashi series introduced in [10]. This connection has been further explored by Zerbes in a series of papers [49, 50, 51], where she related the characteristic element to the generalised Euler characteristic of $X(E/F_{\infty})$. The notion of the generalised Euler characteristic is an extension of the classical Euler characteristic. The idea first appeared in [10] and was later developed and systematically studied by Zerbes in [50].
@@
We now shift our focus to the function field setting, which forms the central theme of this paper. In view of the well-known analogy between number fields and global function fields, one naturally expects that Iwasawa-theoretic phenomena should also manifest in the function field setting. Indeed, several works have explored this direction (see [4, 36, 38, 39, 43, 45, 48]). From now on, let $p$ denote a fixed prime, where we allow $p=2$. Throughout this paper, $F$ will denote a global function field, that is, a finite extension of the field $k(T)$ of rational functions in one variable over a finite field $k$. We always assume that the characteristic $\mathrm{char}(F)$ of $F$ is distinct from our fixed prime $p$. Let $\bar{F}$ be a separable closure of $F$. The cyclotomic ${\mathbb{Z}}_{p}$-extension $F^{\rm cyc}$ of $F$ is obtained by adjoining to $F$ the unique ${\mathbb{Z}}_{p}$-constant field extension of $k$. Writing $\mu_{p^{\infty}}$ for the group of $p$-power roots of unity in an algebraic closure $\bar{k}$ of $k$, the field $F^{\rm cyc}$ is the unique ${\mathbb{Z}}_{p}$-extension of $F$ contained in $F(\mu_{p^{\infty}})$. Following the analogy with the number field setting, we define an admissible $p$-adic Lie extension $F_{\infty}/F$ as a Galois extension satisfying the following properties:
  1. (1)
  2. (2)
  3. (3)
    The extension $F_{\infty}/F$ is unramified outside a finite set of primes.
It is worth noting that a global function field $F$ with $\mathrm{char}(F)\neq p$ admits only one ${\mathbb{Z}}_{p}$-extension, namely the cyclotomic one (see [35, Proposition 10.3.20]). Thus, any admissible extension of dimension at least two is necessarily non-commutative, highlighting the relevance of non-commutative Iwasawa theory in the function field context. In contrast, when $\mathrm{char}(F)=p$, there are infinitely many geometric (i.e., not containing any constant field extension) ${\mathbb{Z}}_{p}$-extensions of $F$ (see [16, Theorem 3]), which marks a significant distinction between the number field and function field cases.
We now return to the main results of this paper. Our first main result is as follows:
Theorem 1.3.
Let $p$ be any prime. Let $A$ be an abelian variety over a global function field $F$ with characteristic prime to $p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then the $\mathfrak{M}_{H}(G)$-conjecture holds for the Pontryagin dual $X(A/F_{\infty})$ of the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$. Furthermore, the Pontryagin dual $X(A/F^{\mathrm{cyc}})$ of the Selmer group ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ has $\mu$-invariant equal to zero, and the generalised $\mu$-invariant of $X(A/F_{\infty})$ is zero. In particular, Mazur’s conjecture holds for $A/F$.
We now make several remarks. This theorem completely resolves Mazur’s conjecture and the $\mathfrak{M}_{H}(G)$-conjecture in the setting of global function fields. This presents a stark contrast to the number field case. It has long been known that the Pontryagin dual of the Selmer group over the cyclotomic ${\mathbb{Z}}_{p}$-extension of a number field can have a positive $\mu$-invariant (see [31, §10, Example 2], [14]). By contrast, the theorem shows that such a phenomenon never occurs in the function field setting when $\mathrm{char}(F)\neq p$. We also note that Sechi [39] previously proved this result for elliptic curves $E$ without complex multiplication, defined over a global function field $F$, where $F_{\infty}$ is obtained by adjoining the $p$-power torsion points of $E$ to $F$, under the assumption that ${\mathrm{Sel}}(E/F)$ is finite.
In the case where $\mathrm{char}(F)=p$, we similarly define $F^{\rm cyc}$ to be the unique constant field ${\mathbb{Z}}_{p}$-extension of $F$. Let $F_{\infty}/F$ be a $p$-adic Lie extension satisfying conditions analogous to (1), (2) and (3) described above. For an abelian variety $A/F$, we assume that all primes of bad reduction are ordinary and lie among the primes that ramify in $F_{\infty}/F$. Under the additional assumption that the $\mu$-invariant of ${\mathrm{Sel}}(A/F^{\rm cyc})$ vanishes, Ochiai and Trihan proved in [36, Theorem 1.9] that the $\mathfrak{M}_{H}(G)$-conjecture holds in this setting. However, recent work by Lai et al. [23] suggests that $\mu$ may be positive when $\mathrm{char}(F)=p$, thereby highlighting a contrast with the $\mathrm{char}(F)\neq p$ case, which is the main focus of this paper.
Our proof of Theorem 1.3 differs entirely from previous works on the ${\mathfrak{M}}_{H}(G)$-conjecture. A key element of our proof is the observation that the cohomological dimension of ${\mathrm{Gal}}(\bar{F}/F^{\mathrm{cyc}})$ is equal to one. From this starting point, our method proceeds via a purely cohomological approach. (Upon completing our work, the authors learned that Witte had proven a similar result on the $\mathfrak{M}_{H}(G)$-conjecture using entirely different methods (see [48]); we thank Jishnu Ray and Fabien Trihan for pointing this out.) While Witte’s result on the $\mathfrak{M}_{H}(G)$-conjecture may initially appear more general, as his approach involves the use of Selmer complexes, our result recovers his, given the well-documented relationship between Selmer groups and Selmer complexes (see [15, Propositions 4.2.35 and 4.3.13]; see also [29, Lemma 3.3.2]). Of course, this same relationship implies that our result follows from that of Witte. That said, we would like to elaborate on the motivation behind the alternative approach adopted in our paper. First of all, the focus of our work differs significantly from that of Witte’s. While Witte’s paper is centered on proving a main conjecture, our study is dedicated to investigating the precise structure of Selmer groups and the order of vanishing of their characteristic elements. More precisely, our direct approach proves that the global and local cohomology groups in the defining sequence of the Selmer group are cofinitely generated over $\Lambda(H)$ (see Lemma 3.3 and the discussion preceding Lemma 5.3), a fact which does not appear to be explicitly established in Witte’s work. In contrast, within the number field context, these cohomology groups are not cofinitely generated over $\Lambda(H)$ (see [37, Theorems 3.2 and 4.1]), necessitating an elaborate snake lemma argument for the computation of the Akashi series of a Selmer group (as in [10, 50, 51]). In the function field context, our observation that these groups are $\Lambda(H)$-cofinitely generated allows for a more straightforward computation of the Akashi series of Selmer groups by directly computing the Akashi series of the global and local cohomology groups (see Section 5 for details). Additionally, we show that the generalised $\mu$-invariant of the dual Selmer group $X(A/F_{\infty})$, as well as the $\mu$-invariant of the dual Selmer group $X(A/F^{\mathrm{cyc}})$, are both equal to zero (see Corollary 3.5). We note that the vanishing of the $\mu$-invariant is already implicit in Witte [48, Corollary 4.21], where it is shown that the dual Selmer group lies in a certain category $\mathbf{N}_{H}$, which essentially also implies that $\mu=0$. Finally, without relying on a non-commutative main conjecture, we present results related to the Birch–Swinnerton-Dyer conjecture, based on Tate’s fundamental work, under an assumption on the finiteness of the Tate–Shafarevich group (see Theorem 1.4 and the subsequent discussion).
In light of Theorem 1.3, we can now attach a characteristic element $\xi_{A}$ to $X(A/F_{\infty})$ in the sense of Coates et al. This enables us to evaluate $\xi_{A}$ at a given Artin representation $\rho:G\to{\mathrm{GL}}_{n}(\mathcal{O})$, where $\mathcal{O}$ is the ring of integers of a finite extension of ${\mathbb{Q}}_{p}$ in $\overline{{\mathbb{Q}}}_{p}$, via the homomorphism $\Phi_{\rho}$ (for details, see Section 7, particularly (7.2)). This yields an element in the field of fractions of the formal power series ring $\mathcal{O}[[T]]$ in an indeterminate $T$ with coefficients in $\mathcal{O}$. To facilitate explicit calculations in the discussion that follows, we introduce the following assumption:
$\mathbf{(G)}:\quad H^{i}(G,A_{p^{\infty}}(F_{\infty}))\text{ is finite for all }i\geq 1.$
We are now in a position to state our next result, presented here in a slightly simplified form. For a more precise andgeneral version, we refer readers to Theorem 7.7, Corollary 7.8, and Theorem 7.12.
Theorem 1.4.
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Denote by $\xi_{A}$ the characteristic element of $X(A/F_{\infty})$, and let $\mathrm{reg}_{F}$ denote the regular representation of $F$. Assuming $(\mathbf{G})$ holds, we have
Moreover, equality holds if Greenberg’s semi-simplicity conjecture (Conjecture 6.2) holds for $A/F$. Additionally, if we assume the finiteness of $\hbox{\cyr X}(A/F)$, then
In fact, by the deep theorem of Kato–Trihan [22], the finiteness of the $\ell$-primary part of the Tate-Shafarevich group $\hbox{\cyr X}(A/F)(\ell)$ for a single prime number $\ell$ is sufficient to establish the final equality in the theorem.
To state our final result, we first introduce the following stronger finiteness assumption. Recall that we write $H={\mathrm{Gal}}(F_{\infty}/F^{\rm cyc})$.
$\mathbf{(H)}:\quad H^{i}(H,A_{p^{\infty}}(F_{\infty}))\text{ is finite for all }i\geq 0.$
We provide a criterion for this condition to hold in Proposition 7.5 of the paper (see also [50, Section 5] and [29, Remark 6.6] for cases where ($\mathbf{H}$) is known to hold in the number field context). Let $S$ be a nonempty finite set of primes of $F$ containing all the primes where the abelian variety $A$ has bad reduction, as well as all the primes of $F$ that ramify in $F_{\infty}$. Let $S^{\prime}$ denote the subset of primes in $S$ whose inertia group in $G$ is infinite. We write $A^{*}$ for the dual abelian variety. The following is our final main result.
Theorem 1.5.
Let $A$ be an abelian variety defined over a global field $F$ with $\mathrm{char}F\neq p$ and let $F_{\infty}$ be an admissible $p$-adic Lie extension of $F$. Assume that $(\mathbf{H})$ holds. Then ${\mathrm{Sel}}(A/F_{\infty})$ has a finite generalised $G$-Euler characteristic $\chi(G,{\mathrm{Sel}}(A/F_{\infty}))$ if and only if ${\mathrm{Sel}}(A/F^{\rm cyc})$ has a finite generalised $\Gamma$-Euler characteristic $\chi(\Gamma,{\mathrm{Sel}}(A/F^{\mathrm{cyc}}))$. Moreover, if this is the case, then
where, for each $v\in S^{\prime}$, $w$ is a fixed prime of $F^{\mathrm{cyc}}$ lying above $v$, and $\Gamma_{w}$ denotes the decomposition subgroup of $\Gamma$ at $w$.
This result is a natural analogue of Zerbes [50], who proved a similar statement in the context of number fields. In the function field context, our theorem generalises earlier computations by Sechi [39] and Valentino [45]. When $A$ is an elliptic curve, our formula recovers Sechi’s result, in which Euler factors of the complex $L$-function appear on the right-hand side. For explicit examples, we refer the reader to [39].
We conclude this introduction with an outline of the paper. Section 2 collects foundational material on Iwasawa algebras, which will be used throughout the paper. The proof of Theorem 1.3 is given in Section 3. The key observation is that the Galois group ${\mathrm{Gal}}(F_{S}/F^{\mathrm{cyc}})$, where $F_{S}$ is the maximal algebraic extension of $F$ contained in $\bar{F}$ which is unramified outside $S$, has $p$-cohomological dimension one (see Lemma 3.2). In Section 4, we establish the surjectivity of the localisation map defining the Selmer group ${\mathrm{Sel}}(A/F_{\infty})$, using Jannsen’s spectral sequence and Nekovář’s duality theorem. This is a key input in our computation of Akashi series in Section 5. In Section 6, we prove a control theorem, showing in particular that the restriction map ${\mathrm{Sel}}(A/F)\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})^{\Gamma}$ has a finite kernel and cokernel. This allows us to relate the order of vanishing at $T=0$ of the characteristic power series of ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ to the ${\mathbb{Z}}_{p}$-corank of the Selmer group ${\mathrm{Sel}}(A/F)$. Section 7 relates the order of vanishing of characteristic elements evaluated at Artin representations to Selmer coranks and their twists over intermediate subextensions of the $p$-adic Lie extension. This yields Theorem 1.4. In Section 8, we introduce the definition of the generalised Euler characteristic, following the framework of Zerbes [50]. Finally, in Section 9, we compute the generalised Euler characteristics under the assumption $(\mathbf{H})$, thereby establishing Theorem 1.5.
2. Prerequisites on Iwasawa algebras
Throughout, let $p$ denote a fixed prime. Let $G$ be a compact $p$-adic Lie group without $p$-torsion. The Iwasawa algebra of $G$ over ${\mathbb{Z}}_{p}$ is defined by
[…]
where $U$ runs over the open normal subgroups of $G$, and the inverse limit is taken with respect to the natural projection maps. Assume for now that $G$ is a pro-$p$ group. Then the ring $\Lambda(G)$ admits a skew field of fractions $\mathcal{Q}(G)$, which allows us to define the rank of a $\Lambda(G)$-module $M$ as follows:
The module $M$ is said to be torsion if ${\mathrm{rank}}_{\Lambda(G)}(M)=0$.
Now, suppose that $G$ is not necessarily pro-$p$. By Lazard’s theorem (see [13, Corollary 8.34]), one can always find an open subgroup $G^{\prime}$ of $G$ that is pro-$p$. In this case, a $\Lambda(G)$-module $M$ is said to be torsion over $\Lambda(G)$ if it is torsion over $\Lambda(G^{\prime})$ in the sense defined above. We shall also make use of the following equivalent characterisation of a torsion $\Lambda(G)$-module: ${\mathrm{Hom}}_{\Lambda(G)}(M,\Lambda(G))=0$ (cf. [46, Remark 3.7]).
3. Selmer groups and the ${\mathfrak{M}}_{H}(G)$-Conjecture
Let $p$ be any prime number. Throughout, let $F$ denote a global function field of characteristic coprime to $p$, and write $k$ for its constant field. Suppose that $A$ is an abelian variety defined over $F$. Let $F_{\infty}$ be a $p$-adic Lie extension of $F$ satisfying the following admissibility conditions: